using the method of undetermined coefficient find the general solution of the differential equation dy^4/dx^4-2dy^3/dx^3+2dy^2/dx^2=3/e^x+2x/e^x +sinx /e^x
1
Expert's answer
2020-05-19T13:42:13-0400

this equation can be written as (D4-2D3+2D2)y = (3+2x)e-x+sinx.e-x ....(1)

the auxiliary equation is

D4-2D3+2D2=0

D= 0,0,1+i,1-i

so, the complimenatry function is

y= c1+c2x + ex(c3 cosx + c4sinx)

let trial solution be

z= e-x(A sinx + B cosx) + e-x(D+Ex)

Dz= -e-x( A sinx + B cosx + D+Ex) + e-x( A cosx - B sinx + E )

D2z= e-x( D+Ex) -2e-x(A cosx - B sinx +E)

D3z = -e-x(-2A sinx -2B cosx + D+Ex) + 3e-x(A cosx - B sinx +E)

D4z = e-x( -5A sinx -5 B cosx + D+Ex) - e-x(A cosx - B sinx + 4E )


putting these values in (1), we get

e-x[ sinx(-9A+9B) + x (5E) + 5D -14E + cosx(-9B-9A)] = e-x(3+2x+ sinx)

comparing coefficients


A=-1/18

B=1/18

D=43/25

E= 2/5


So, the solution is

y= c1+(c2+0.4e-x)x + ex(c3 cosx + c4sinx) + e-x[( -1/18) . sinx + (1/18) . cosx + (43/25 )]